A jogger runs 186 meters westward, then 210 meters eastward, then 156 meters westward, and finally 84 meters eastward.
For this motion, what is the distance moved?
What is the magnitude and direction of the displacement?
1
Expert's answer
2020-05-21T13:45:40-0400

The distance moved is the sum of all distances the jogger traveled: 186+210+156+84=636 m.

The magnitude of the displacement is the distance between the initial and final positions of the jogger. Assume that distances westward are negative (red) and they are positive (green) when the jogger runs eastward:


"d=-186+210-156+84=-48\text{ m}."

Since the displacement is negative, the direction of the displacement is westward: